OpenStudy (anonymous):

Simplify: 2m^2⋅4m^3/2⋅4m−^2 ONLY THE SMART WILL UNDERSTAND!

OpenStudy (anonymous):

@hayonce123 u rite only the smart will understand cuz i dont nun of this pellet

OpenStudy (agreene):

which is it? \[\frac{2m^2*4m^3}{2*4m^{-1}}\large\space\text{or}\space 2m^2*4m^{\frac32}*4m^{-2}\]

OpenStudy (agreene):

I'm gonna assume its the first. \[\frac{2m^2*4m^3}{2*4m^{-2}}=\frac{8m^8}{8m^{-2}}=m^7\]